OpenStudy (anonymous):

Screenshots! Already Answered! Circle questions, just need someone to check the both of them!

OpenStudy (nubeer):

hmm i think first one is wrong.

OpenStudy (nubeer):

second looks right.

OpenStudy (hba):

The other one looks good to me. The first one is wrong \[(x-a)^2+(y-b)^2=r^2\]

OpenStudy (slaaibak):

First one should be C.

OpenStudy (slaaibak):

shot bro.

OpenStudy (hba):

As can be seen from your equation \[(x-3)^2+(y-2)=4^2\] So the term does not have a square on it .

OpenStudy (anonymous):

first one is not a circle because there is no \(y^2\) term, pick C